A trip to New York City April 2003

Edition 1, Version 1

by Jorgen Skogstad and Santa Claus

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2003 Jorgen Skogstad, email: jskogsta@pogostick.net

Introduction

Well, this trip was planned quite sometime ago, but here we are; New York city. We’ve been here for some days already and still the impressions of this metropolis have not sunk in yet. There is too much to comprehend, understand and gather about this city. Far too much to get a grip over in just the few days we’ve been here. However; one thing is for sure, New York city gets to you.

2 A trip to New York City, april 2003.

1 Preparations for New York

When you are going to New York you have to make certain preparations. Here are some easy guidelines worth looking through before you go.

Stuff to remeber upon arrival in New York;

One dollar bills for tipping

If it is your first trip to New York be prepared for the initial cab ride and the porter at the hotel. They expect to be tipped and dragging out a 20 dollar bill is a bit extravagant. You’re not (I presume!) staying at the Walldorf! Give the cab driver 3-8 dollars depending on the extent of the drive. The porter would presumably expect 2-5 dollars for bringing up you luggage and telling you the workings of the hotel.

Travel adapter for electrical conversions

If you are bringing your handicam and digital camera do remember to bring your travel adapter for european (if you are from there!) to american electrical conversions. Also bear in mind that the american system is working at 110/120 volts, not the average 220 volts we have most places in Europe. I forgot bying the one on offer inflight at Air France which was too bad. I really could not find any decent worldwide electrical adapters here in New York. Understandably they most offer the american to the world converters, not the other way around.

Good shoes!

If you plan on getting around you end up walking alot. Therefore bring your sneakers. They are essential in any visit to New York!

Rucksack You are bound to do shopping here in New York. More than you ever think you are going to. Therefore; bring a good rucksack to haul your stuff around. It will save you alot of grief!

Get the Metrocad (7 day one??!)

It is an essential must have for any visitor in New York. Of course you could take the cab all the time, but you end up spending alot of money and they are not always that much faster. With the 7 day Metrocard (we got) you will be able to use the subway and the busses for unlimited rides within the period. Good value at 17 dollars! Go get it at first chance!

4 A trip to New York City, april 2003.

2 Planning the trip

My father, Erik, have always had an interest in family relations and kinship. The ‘basis‘ for this trip was founded in this. He wanted to visit New York, since this was for most european migrants, their first stop to the promised lands of the Americas during the early stages of this century. The primary goal was Ellis Island (See Appendix A [References], page 13) where we would look through their registers for clues of some of our ancestors that had migrated to the states.

During february we started the initial planning for the trip. Initially it was only myself and my father who was planning to go, but my sister Merete also joined in. She decided to bring her eldest son, Trygve as well. We ended up being the four of us.

More!!!

6 A trip to New York City, april 2003.

3 Arriving in New York city

The airplane speakers burst out ‘We are approaching Newark international airport. Blah blah blah ... immigration cards and information for customs personell needs to be filled out upon arrival ... blah blah .. ‘. The thoughts on what to come speed up. Images and impressions of the past surfaces and anticipations of what to come increases.

New York have always been portrayed as the megacity of all cities. It is by all standards the definition of urbanity. At least when considering Manhatten with it’s skyline that everyone have at least seen a dozen times on the movie- and tv-screen. As such, I guess everyone visiting this city has expectations on what this city will actually bring. So did I (and we!)!

Getting off the airplane and through customs and immigration was a breeze. Standing outside the airport gives a special sense of completion. We’ve at least had a day of travel behind us and the goal is just around the corner. We can even see the Manhatten skyline from the airport. Now the only thing behind us and downtown Manhatten is the twenty minute drive with a yellow NY taxicab.

It was a breeze getting the taxi, just standing in queue waiting for our turn. The cab driver we got must have been a jamaican who must have lived here for quite some time. As the fables tell you, in New York you stick to yourself. That was so with the cabdriver we had. Even though I tried to pick up a conversation about the marvelous weather we had he said a few gibberish phrases I could hardly understand and that was it. He was done. I wondered if this was to be typical behaviour of the New Yorkers I were to encounter later (and to some degree it was!).

Driving towards the inner city holds a certain amount of anticipation even though I’ve been around. Even cities like San Francisco, London, Toronto and Paris cannot match the awe giving sight of the New York city skyline (yes; even without the twin towers!). Driving down the highway and under the Hudson river builds the expectations of getting into the grid of skyscrapers Manhatten consists of. And getting through the tunnel under the Hudson and into Manhatten itself brings a strange sense of awe! You end up watching the people, cars and buildings passing by. Even seeing the top of the buildings from the car was hard. After a few minutes the cab stops beside our hotel, The Grand Hyatt New York (See Appendix A [References], page 13), and drops us off. Not having enough short change for tip I sense the cabbie’s disapointment. ’Only two dollars tips?’. Well dude, you should have brought up a good conversation and we could have perhaps managed something more? Well, anyways; we’ll never meet again.

The hotel was pretty impressive. A 34 story hotel just beside (on top?) the Grand Central station on 42nd street. Hmm, searching the web and ordering through Priceline (See Appendix A [References], page 13) proved worthwhile. We got two doubles on the 25th floor.

8 A trip to New York City, april 2003.

4 Recomendations and tips of things to do in New York city

With the sheer size of New York city you could spend years trying to cope with the offerings within this city. No way you would be able to get through it all. One thing is for sure; you ought to sit down and really think about what you want to do with your stay here. If you have just a few days, like us, this is vital to get the most out of your stay. We had like 8 days here and still only scratched the surface of the offerings.

4.1 Sights and things to do

There are alot of sights available in New York and Manhatten. Some worth seeing, others not. Well, here I present some that we found either good or not and a description of each one.

Empire State Building

After the destruction of the World Trade Center this have become the highest skyscraper in New York. Well worth a visit. On the visitin platform there is an awsome view of the city. One peculiar thing to bear in mind when one is up there is that two Norwegian skydivers actually did a basejump from this building some years ago. The story goes that they tricked the security guards and then managed to climb the railings and go overboard. They landed without problems and huddled their stuff together and proceded down into the subway before police could respond! Cool!

New York Skyride

This is actually a combined venue within the Empire State building. It is a virtual heliride over downtown New York and then some, but for some reason we thought it to be a bit ‘too much and a tad American‘. We were actually a bit disapointed with it. As always Americans are really good at marketing, but when it comes to proving the ‘goods‘ they sometimes fails. This is true with the New York skyride. Well; I can’t say that I recomend this. Caution be advised and don’t come running to me if you were disapointed as well!

Ellis Island

The boat to Ellis Island and the statue of liberty goes from Battery Park. It takes some 15-20 minutes out to the first stop at the statue of liberty before proceeding to Ellis Island. Ellis Island is an interesting site where most of the us immigrants met the promised lands. (There should be more inforation here!)

Statue of Liberty

The statue of liberty is a must see when you are in New York. Not that is extremely interesting, but it is the embodyment of the American spirit. This is America! There are in fact a few interesting points about the statue of liberty;

1. The Statue of Liberty was a gift from France!

2. The original design copy/test is placed by the river Seine in Paris, France!

3. Most of the copper on the statue is from Karmoey in Norway!

Grand Central Park

A big site! You could very well spend a whole day or several here. If you are staying here for several days the park is a good place to cool off between your days of sightseeing and shopping.

Movie and theatre tickets at Times Square

Times Square is well worth a visit. Especially at evenings when the lights are out and it is at is busiest. Movie tickets are typically 10 dollars per person. Not sure about the theater tickets since we decided not to go. However you could get discounted theater tickets just off Times Square, but beware of loooong queues. If you are serious about getting these discounted tickets you should get the opening hours of TKTS just off Times Square and hunch out there on hour or so before they open. Just to beat most of the queue and have a fair chance to getting some decent ones.

4.2 Recomended shops

Macy's Suposedly the worlds largest store. Was actually a bit disapointing when we were there. Seemed a bit run down, but still the available goods there is just plain overwhelming. Would be nice to have been there during their sale season!

Toys'r'us

Several shops around the city (and the world!). The largest is by far in the Times Square area. Suposedly 1 million toys. When we were there it had an inhouse ‘paris wheel‘ over fours stories. Impressive! If you want to have your child yelling for toys and being really annoying this is the place for you.

FAO Schwarz

Situated within the Trump Tower on 5th avenue. Impressive store for the quality consious. Even though I was not looking for toys I found this really interesting and impressive.

Manhatten portage

A small shop selling durable portables. If you are looking for a bag of some sort that will endure close to anything this might be the place for you. Prices seem fair as well. Most of the products are nylon based and come in almost every colour available.

J&R computers (or something similar!)

This family owned store has become a one stop shopping block for all your electronic needs. They carry photo (digital and analog), computer, games, music and the lot at decent prices. They are situated just off the Wall Street area at XXth and XXnd.

Centry 21 This department store is a must for the fashion oriented male, female and child. You can find designer brands here at low prices. The prices are typically 30-50% off the American prices of the brands which again are a few percent lower than the ones we have in Norway. There are numerous outlets throughout the city, but presumably the biggest one is just off the World Trade center site. You can’t miss it if you go to the World Trade center site.

5 On the way home!

Well, the time had come for the return to the homeland! Well, as always the return comes with a certain anticipation. Even though it always is fun going to new places and doing new things home is always home. The place for relaxation and just plain family life!

We got a limosine service out of Manhatten. What other way to leave the ‘promised land‘ than in the defacto American limo? With a spanish accented guy driving the biggest flake I’ve driven in for a long time. Watching the skyline depart in the rear mirror. Wierd feeling, but still .. it feels good. Being on the return to home.

What struck me when we got to the airport is the feeling that Americans have no idea on organizing stuff. They seem to think that throwing an enormous amount of people at things always will solve the problem. Jeez, they are wrong! For the duration of our stay I must say that I am not impressed with the organising skills. They are really good at forming queues, but how to organise their disipation they have no clue! I always end up (for some strange reason?!) thinking that ‘have we just been in Germany, THEN things would have moved along alot faster‘. Then again Germans are known for their ‘Ordnung muss sein¡.

Thinking back it seems that you will find double the amount of people in any American department store or shop than you would in the average european counterpart. How can that be? Is this just a governmental plot from the Americans to feed their economy and provide a certain amount of employment? I would be inclined to believe so. Even in the smallest of food stalls you will most always find two to three people serving you. Why is this? Is this even feasible and economically viable? My instant guess would be no, but then again I know little to nothing about the working of American small business ownership and operations. Just strikes me as a bit odd.

Anyways; when you are about to enter a trans continental flight you always end up wondering what kind of people you will sit beside. I remember a flight to San Fransisco (from London) in which I was placed right in the corner squeezed in between what must have been 7-9 indians og pakistanis. I have nothing against foreign cultures and so, but that flight was a drag. Pure hellish to say the least. Boy were they nosiy! Not to mention the two small children they had. They were on top of everyone, even me who was not part of their party. It seemed that they did not bother about little me and threw things about most certain to know me over the few times I managed to close my eyes. Hmmm. So, what about this flight? Newark to Paris should be a safe trip, but what must be czhechs or russian youths kept half of our section busy. No respect for others. Noise and the likes. Hmm, I wonder why I always end up with the short straw on these long haul flights! Well, there must be a plot I am not aware of.

Anyways, the clock not is about 08:31 and we have landed a couple of hours ago at Charles de Gualle airport in Paris. Just wainting for our connecting flight to Gothenburg in Sweden. So far so good. Sitting sipping a strong coffee and typing these very words. A good start to a new day .. which most certainly have to bring a donut or something soon since I am about to starve!

12 A trip to New York City, april 2003.

Appendix A References

This appendix lists some interesting references for this trip. These references were working at the initial revisions of this document (if they for some reason have stopped working by the time you read this!).

Priceline

Priceline is a US based website offering reduced rates on travel items like hotels, rental cars etc. The neat thing about Priceline is that it is auction based. The registered users put forth an offer based on a set of search criteria and if they’re in luck; some hotel (or others!) accept the offer and they’re set to go! We managed to get a 4 star hotel room for $150, when the registered price was $349. We could possibly push the price even futher down, but still it wasn’t too bad based on the location and the quality of the hotel. Priceline

(http://www.priceline.com/).
Grand Hyatt New York Hotel

The hotel we stayed in located centrally at the Grand Central station on 42nd

street. Grand Hyatt New York (http://www.grandhyattnewyork.com/).

Mr.Jet Mr.Jet is a swedish website offering online travel services. Through Mr.Jet we managed to get return fare tickets from Gothenburg Sweden to Newark NY for aproximately 2500 norwegian crowns. That’s pretty cheap!

http://www.mrjet.com/, Mr.Jet.
Jorgen Skogstad's websites

I have several websites and domains. The ones I could remember I have listed beneath.

 1. http://www.pogostick.net/~jskogsta/, jskogsta@pogostick
 2. http://www.skogstad.no/, skogstad.no
 3. http://www.skogstad.com/, skogstad.com
 4. http://www.nobliege.net/, nobliege.net
 5. http://www.gintonics.com/, gintonics.com
 6. http://www.helpdesken.no/, helpdesken.no
 7. http://www.securitywarning.com/, securitywarning.com
Ellis Island

Ellis Island was for most european migrants their first stop on their route to the promised lands during the early stages of this century.

http://www.ellisisland.com/, Ellis Island.

14 A trip to New York City, april 2003.

Concept Index 15

Concept Index

I  ^{Intex entry for first chapter . . . . . . . . . . . . . . . . . . . . 1}

16 A trip to New York City, april 2003.

Table of Contents

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Preparations for New York . . . . . . . . . . . . . . . . . 3 2 Planning the trip . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3 Arriving in New York city. . . . . . . . . . . . . . . . . . 7 4 Recomendations and tips of things to do in

New York city . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

4.1 Sights and things to do . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.2 Recomended shops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

5 On the way home! . . . . . . . . . . . . . . . . . . . . . . . . 11 Appendix A References . . . . . . . . . . . . . . . . . . . . . 13 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

ii A trip to New York City, april 2003.